OpenStudy (anonymous):

Find the nth term of the arithmetic sequence where a1 = 5000 and d = -100

OpenStudy (muhammad_nauman_umair):

a (nth)=a(0ne)+(n-1)d

OpenStudy (muhammad_nauman_umair):

a (nth)=5000+(n-1)(-100) a (nth)=5000-100n+100 a (nth)=5100-100n